# SAGe: A Lightweight Algorithm-Architecture Co-Design for Alleviating Data Preparation Overheads in Large-Scale Genome Analysis

### v1

@@ 7.2. Area and Power
Table 2 lists the area and power consumption of SAGe’s accelerator units operating at 1 GHz. Although these units could operate at a higher frequency, their throughput is already sufficient because SAGe’s accelerator operations are bottlenecked by the NAND flash read throughput. SAGe’s accelerators consume a small area and power of 0.0023 mm2 and 0.95 mW at 22 nm node. The area overhead of SAGe hardware is only 0.7% of the three 28-nm ARM Cortex R4 cores (cortexr4, ) in a simple SSD controller (samsung860pro, ).  [^18] For a conservative comparison, we scaled SAGe’s area to 32nm node using (stillmaker2017Scaling, ).
@@ 7.3. Energy
We evaluate energy consumption for the end-to-end process of preparing data (reading data from the storage system, and decompressing and reformatting the read data) and performing read mapping on it. Read mapping is a fundamental process in the vast majority of genome analysis use cases (li2018minimap2, ; alser2020technology, ; cali2020genasm, ), for which it has been an optimization target in a large body of prior work (zhang2000greedy, ; slater2005automated, ; li2018minimap2, ; myers1999fast, ; marco2021fast, ; marcosola2023optimal, ; grootkoerkamp2024apa2, ). SAGe reduces the end-to-end energy by alleviating the energy overhead of preparing data for the energy-efficient read mapping accelerator. We evaluate the energy consumption based on the energy of the host processor and DRAM, accelerator units, SSD operations, and host/SSD/accelerator communication. Each part’s energy is based on its active/idle power and execution time. Fig. 16 shows the end-to-end energy reduction over pigz. We observe that, due to its lightweight structures and efficient access patterns, SG leads to a significant average energy reduction of 49.6$\times$, 24.6$\times$, and 18.8$\times$, over pigz, (N)Spr, and (N)SprAC, respectively.
@@ 7.4. Compression Ratio
Table 3 shows the compression ratios of pigz, (N)Spr, and SG. We observe that SG achieves (i) 2.9$\times$ better average compression ratio than pigz, and (ii) comparable compression ratios to the state-of-the-art genomic compression, (N)Spr, (with a modest 4.6% average reduction). SAGe achieves these high compression ratios while providing fast decompression speed. SAGe software (hardware) achieves 11.6$\times$ (53.0$\times$) and 3.3$\times$ (15.1$\times$) average decompression speedup over pigz  and (N)Spr, respectively. As shown in §7.1, this faster decompression leads to substantial improvement in the end-to-end throughput of genome analysis pipelines.
@@
To show the impact of different optimizations, Fig. 17 presents the size breakdown of the reads’ mismatch information [^19] Rev refers to a bit marking if a read matches in reverse to the consensus sequence. for a short (RS2) and long read set (RS4) in five settings: (iNO, with no optimization on the raw mismatch information; (iiO1, with matching position optimization (§5.1.3) added; (iiiO2, with mismatch positions and count optimizations (§5.1.1) added; (ivO3, with mismatching base and type optimizations (§5.1.2) added; and (vO4, with corner case optimizations (§5.1.4) added.
@@ 8. Related Work
Accelerating Genome Analysis. Many prior works accelerate genome analysis via algorithmic optimizations (e.g., (zhang2000greedy, ; slater2005automated, ; li2018minimap2, ; myers1999fast, ; marco2021fast, ; marcosola2023optimal, ; grootkoerkamp2024apa2, )), or hardware acceleration using GPUs (e.g., (jia2011metabing, ; kobus2021metacache, ; wang2023gpmeta, ; kobus2017accelerating, ; Su2012, ; su2013gpumetastorms, ; Yano2014, ; Gamaarachchi2020_f5c, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; wilton2015arioc, ; liu2009cudasw++, ; liu2010cudasw++, )), FPGAs (e.g., (saavedra2020mining, ; zhang2023genomix, ; cervi2022metagenomic, ; shih_haru_2023, ; fujiki2020seedex, ; banerjee2018asap, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ; liyanage2023efficient, )), ASICs (e.g., (turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, )), processing-in-memory (e.g., (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, )), and ISP (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )). There are also many accelerators for specific parts of the genome analysis pipeline  (saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; fujiki2020seedex, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ). SAGe is orthogonal to these works and can be flexibly integrated with them to further improve their benefits.
@@
Genomics-Specific Compression. Due to inefficiencies of general-purpose compression for genomic data, a large body of works (e.g., (chandak2018spring, ; Deorowicz2020, ; lan2021genozip, ; alyami2019lfastqc, ; kowalski2019pgrc, ; roguski2018fastore, ; chandak2017compression, ; cogo2021genodedup, ; Meng2023, ; kokot2022colord, ; dufort2020enano, ; dufort2021renano, ; karasikov2022lossless, ; vandamme2024tinted, ; dragenora, )) improves compression ratios via genomics-specific compression. As detailed in §4, some works (e.g., (guo2013gpu, ; qiao2019fpga, ; zhao2017streaming, ; jiang2021exma, ; arram2015fpga, ; wang2018accelerating, ; chen2023efficient, ; leavline2013hardware, )) accelerate certain computational kernels that are widely used in state-of-the-art genomic compressors. Despite their benefits, these works have two key limitations. First, as detailed in §4, they are unsuitable for implementation in the resource-constraint environment inside modern SSDs. Second, they do not fully alleviate end-to-end data preparation overhead for genomic accelerators (detailed in §4 and evaluated in §7). While some works (rajarajeswari2011dnabit, ; saada2016dna, ) use algorithms that do not rely on expensive units, they achieve poor compression ratios (§4 ).
@@
General-Purpose Compression. Many works propose general-purpose compression in software (e.g., (collet2018zstandard, ; pavlov20167, ; Brotli, ; oswal2016deflate, ; goyal2021dzip, ; goyal2018deepzip, ; chen2024ha, )) or hardware (e.g., (bartik2015lz4, ; liu2018data, ; fowers2015scalable, ; chen2021fpga, ; angerd2022gbdi, ; gao2024beezip, ; karandikar2023cdpu, ; 9499902, ; abali2020data, )). However, they achieve poor compression ratios for genomic data compared to genomic compressors (hernaez2019genomic, ; zhu2013high, ). For example, the compression accelerators Intel QAT (intelqat, ) and IBM zEDC (ibmzedc, ) lead to 2.6×2.6$\times$2.6 × and 2.7×2.7$\times$2.7 × lower average compression ratios, respectively, than genomic compressors on our evaluated datasets. Due to their limitations in leveraging longer-range similarities common in DNA data (hernaez2019genomic, ; zhu2013high, ), even state-of-the-art strong general-purpose compression algorithms do not achieve the compression ratios on par with genomic compression tools. For example, we evaluate xz (xz, ) (a state-of-the-art LZMA-based compression tool) and zstd (collet2018zstandard, ) (a recently-developed compression algorithm), both with their highest compression levels, and show that they achieve 2.13$\times$ lower average compression ratios than the state-of-the-art genomic compressors across our genomic datasets. By leveraging the unique features of genomic data, SAGe achieves compression ratios similar to the state-of-the-art genomic compression (with a modest reduction of 4.7%, as shown in §7.4), while also providing faster decompression. Our evaluations show that the software implementation of SAGe achieves 2.1$\times$ and 3.9$\times$ average speedup over xz and zstd, respectively.
In-Storage Processing. Many prior works propose ISP architectures for various applications (liang2019cognitive, ; kim2020reducing, ; lim2021lsm, ; li2021glist, ; wang2016ssd, ; lee2020neuromorphic, ; kang2021s, ; han2021flash, ; wang2022memcore, ; wang2018three, ; han2019novel, ; choi2020flash, ; mailthody2019deepstore, ; pei2019registor, ; jun2018grafboost, ; do2013query, ; seshadri2014willow, ; kim2016storage, ; riedel2001active, ; riedel1998active, ; lee2022smartsage, ; jeong2019react, ; jun2016storage, ; li2023ecssd, ; wang2024beacongnn, ; jang2024smart, ; lee2024presto, ; mahapatra2024instoragedomainspecificaccelerationserverless, ; 10.1145/3689336, ; 10.1145/3632951, ; 10.1145/3631882.3631896, ), such as machine learning (e.g., (li2023ecssd, ; liang2019ins, ; lee2022smartsage, ; wang2024beacongnn, ; jang2024smart, ; pan2024instinferinstorageattentionoffloading, )), genome analysis (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )), and graph analysis (e.g., (jun2018grafboost, )). Some works integrate additional processing units into the SSD controller (gu2016biscuit, ; kang2013enabeling, ; wang-eurosys-2019, ; acharya-asplos-1998, ; keeton-sigmod-1998, ; mailthody2019deepstore, ; pei2019registor, ; jun2018grafboost, ; do2013query, ; seshadri2014willow, ; kim2016storage, ), utilize flash cells (park2022flash, ; choi2020flash, ; han2019novel, ; shim2022gp3d, ; merrikh2017high, ; tseng2020memory, ; wang2018three, ; lue2019optimal, ; parabit_2021_gao, ; yu2024cambriconllmchipletbasedhybridarchitecture, ; ares2024, ), or tightly integrate SSDs with FPGAs (koo-micro-2017, ; jun2015bluedbm, ; torabzadehkashi2019catalina, ; lee2020smartssd, ; ajdari2019cidr, ; 10.1145/3649329.3658265, ) or GPUs (cho2013xsd, ). SAGe is orthogonal to prior ISP genomics works and enables their adoption without relying on large data to be stored uncompressed.
@@ 9. Conclusion
We propose SAGe, an algorithm-system co-design for capacity-efficient storage and high-performance access of large-scale genomic data. We leverage the features of genomic data in compression mechanism and storage architecture design. SAGe can be seamlessly integrated with a broad range of genome analysis systems. SAGe significantly improves the end-to-end performance and energy efficiency of genome analysis accelerators, while achieving high compression ratios and at a low hardware cost.

### v2

@@ 8.2. Area and Power
Table 2 lists the area and power consumption of SAGe’s accelerator units operating at 1 GHz. Although these units could operate at a higher frequency, their throughput is already sufficient because SAGe’s accelerator operations are bottlenecked by the NAND flash read throughput. SAGe’s accelerators consume a small area and power of 0.0023 mm2 and 0.95 mW at 22 nm node. This small overhead enables SAGe to unobtrusively integrate with genome analysis systems even in resource-constrained environments. For example, as shown in §8.1, SAGe can provide significant benefits by enabling integration with NDP accelerators inside an SSD, while consuming only 0.7% of the area of the three 28-nm ARM Cortex R4 cores (cortexr4, ) in a simple SSD controller (samsung860pro, ).  [^14] For a conservative comparison, we scaled SAGe’s area to 32nm node using (stillmaker2017Scaling, ).
@@ 8.3. Energy
We evaluate the energy consumption based on the energy of the host processor and DRAM, accelerator units, SSD operations, and host/SSD/accelerator communication. Each part’s energy is based on its active/idle power and execution time. Fig. 16 shows the end-to-end energy reduction over pigz. We observe that, due to its lightweight structures and efficient access patterns, SG leads to a significant average energy reduction of 49.6$\times$, 24.6$\times$, and 18.8$\times$, over pigz, (N)Spr, and (N)SprAC, respectively.
@@ 8.4. Compression Ratio
Table 3 shows the compression ratios of pigz [^15] We compare against a wider set of general-purpose compressors in §9., (N)Spr, and SG. SG achieves (i) 2.9$\times$ better average compression ratio than pigz, and (ii) comparable compression ratios to the state-of-the-art genomic compression, (N)Spr, with a modest 4.6% average reduction.
@@
To show the impact of different optimizations, Fig. 17 presents the size breakdown of the reads’ mismatch information [^16] Rev refers to a bit marking if a read matches in reverse to the consensus sequence. for a short (RS2) and long read set (RS4) in five settings: (iNO, with no optimization on the raw mismatch information; (iiO1, with matching position optimization (§5.1.3) added; (iiiO2, with mismatch positions and count optimizations (§5.1.1) added; (ivO3, with mismatching base and type optimizations (§5.1.2) added; and (vO4, with corner case optimizations (§5.1.4) added. We make six observations. First, O1 significantly reduces the data size of the matching positions in short reads. Second, O2 significantly reduces the data size of mismatch counts in short reads. Third, O2 leads to a large reduction in the mismatch positions’ data size in long reads. Fourth, O3 significantly reduces the bases for long reads by efficiently encoding chimeric reads. However, the data size for mismatch positions increases in O3 due to the additional mismatches at the chimeric reads’ new matching positions. This tradeoff is suitable for SAGe, as it can efficiently encode the greater number of mismatch positions. Fifth, O3 reduces the types’ data size for short and long reads. Sixth, O4 reduces the size required for labeling corner cases in short reads.
@@ 9. Related Work
To our knowledge, SAGe is the first algorithm-system to mitigate the data preparation bottleneck in a wide range of genome analysis hardware accelerators.
@@
Accelerating Genome Analysis. Many prior works accelerate genome analysis via algorithmic optimizations (e.g., (zhang2000greedy, ; slater2005automated, ; li2018minimap2, ; myers1999fast, ; marco2021fast, ; marcosola2023optimal, ; grootkoerkamp2024apa2, )), or hardware acceleration using GPUs (e.g., (jia2011metabing, ; kobus2021metacache, ; wang2023gpmeta, ; kobus2017accelerating, ; Su2012, ; su2013gpumetastorms, ; Yano2014, ; Gamaarachchi2020_f5c, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; wilton2015arioc, ; liu2009cudasw++, ; liu2010cudasw++, )), FPGAs (e.g., (saavedra2020mining, ; zhang2023genomix, ; cervi2022metagenomic, ; shih_haru_2023, ; fujiki2020seedex, ; banerjee2018asap, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ; liyanage2023efficient, )), ASICs (e.g., (turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, )), and NDP (e.g., (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ; mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )). SAGe is orthogonal to these works and can be flexibly integrated with them to further improve their benefits.
@@
Genomics-Specific Compression. A large body of works (e.g., (chandak2018spring, ; Deorowicz2020, ; lan2021genozip, ; alyami2019lfastqc, ; kowalski2019pgrc, ; roguski2018fastore, ; chandak2017compression, ; cogo2021genodedup, ; Meng2023, ; kokot2022colord, ; dufort2020enano, ; dufort2021renano, ; karasikov2022lossless, ; vandamme2024tinted, ; dragenora, )) improves compression ratios via genomic compression. As detailed in §3.3, some works (e.g., (guo2013gpu, ; qiao2019fpga, ; zhao2017streaming, ; jiang2021exma, ; arram2015fpga, ; wang2018accelerating, ; lim2025bancroft, ; chen2023efficient, ; leavline2013hardware, )) accelerate certain computational kernels that are widely used in state-of-the-art genomic compressors. Despite their benefits, these works (i) are unsuitable for implementation in resource-constrained environments and/or (ii) do not fully mitigate the end-to-end data preparation bottleneck (evaluated in §8). While some works (rajarajeswari2011dnabit, ; saada2016dna, ) use algorithms that do not rely on expensive resources, they achieve poor compression ratios (evaluated in §3.3 ). As a strongly conservative evaluation, 0TimeDec (in Fig. 13) can serve as an idealized representation of decompressors that, despite their optimized performance, cannot be integrated in resource-constrained environments. As shown in §8.1, SAGe’s ability to integrate with accelerators under constrained resources leads to significant benefits over 0TimeDec.
@@
General-Purpose Compression. Many works propose general-purpose compression in software (e.g., (collet2018zstandard, ; pavlov20167, ; Brotli, ; oswal2016deflate, ; goyal2021dzip, ; goyal2018deepzip, ; chen2024ha, )) or hardware (e.g., (bartik2015lz4, ; liu2018data, ; fowers2015scalable, ; chen2021fpga, ; angerd2022gbdi, ; gao2024beezip, ; karandikar2023cdpu, ; 9499902, ; abali2020data, )). However, they achieve poor compression ratios for genomic data (hernaez2019genomic, ; zhu2013high, ). For example, the compression accelerators Intel QAT (intelqat, ) and IBM zEDC (ibmzedc, ) lead to 2.6×2.6$\times$2.6 × and 2.7×2.7$\times$2.7 × lower average compression ratios, respectively, than genomic compressors on our datasets. Due to their limitations in leveraging longer-range similarities common in DNA data (hernaez2019genomic, ; zhu2013high, ), even state-of-the-art strong general compression algorithms do not achieve the compression ratios on par with genomic compression. For example, xz (xz, ) (a state-of-the-art LZMA-based compressor) and zstd (collet2018zstandard, ), both at their highest levels, achieve 2.13$\times$ lower average compression ratios than the state-of-the-art genomic compressors on our datasets.
@@ 10. Conclusion
We propose SAGe, an algorithm-system co-design for capacity-efficient storage and high-performance access of genomic data, to mitigate the data preparation bottleneck. SAGe can seamlessly integrate with a wide range of genome analysis systems. We leverage the features of genomic data in the compression mechanism and architecture design. SAGe significantly improves the end-to-end performance and energy efficiency of genome analysis, while achieving high compression ratios at a low hardware cost.